There was a major collision of an asteroid with the Moon in medieval times. It was described by monks at Canterbury Cathedral in England as a red glow on and around the Moon. How long after the asteroid hit the Moon, which is 8.71 × 105 km away, would the light first arrive on Earth?


2 decimal places


1
Expert's answer
2021-05-30T13:23:34-0400

"t=\frac sc,"

"t=\frac{8.71\cdot 10^5}{3\cdot 10^5}=2.9~\text s."